# The syzygy distinguisher

### v4

@@ 4 Regularity $2$ and the small defect heuristic
Generating series reformulation of [9, Cor. 1.10].
Definition 5 ([25, Def. 1.5 & Th. 2.35])
@@
Definition 6 (cf. [9, §4A], after [22, Lect. 14])
@@
$\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}.$ (83)
@@
Regularity $2$ helps in computing the full Betti diagram. We already gave examples in Figures 3-3. In order to provide a few more, let us first recall that a $[n,k]$ code ${\mathsf{C}}$ is MDS if it has dual minimum distance $\operatorname{d}_{\min}({\mathsf{C}}^{\perp})=k+1$.
Lemma 9 ([14, Th. 1], reformulated)
Proposition 8
$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (85)
Proof
Proposition 9
  • $\beta_{0,0}=1$
  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
Proof
Now set $n=2k-1$.
Figures 5 and 5 illustrate Propositions 8 and 9.
Figure 4: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
Figure 5: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
Exercise 1
@@
A direct application of Definition 7 and Theorem 4.1 gives:
Proposition 10
If ${\mathsf{C}}$ is a $[n,k]$-code with ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$, then for all $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r})$ (86)
$\beta_{r-2,r}({\mathsf{C}})=\left(\frac{k(k+1)}{r}-n\right)^{-}\binom{k-1}{r-2}+\operatorname{def}(\varphi_{r}).$ (87)
It is tempting to see this result as an estimate for the Betti numbers of random codes, with a leading term given by $\operatorname{ind}(\varphi_{r})^{\pm}$ (compare with the bound in Lemma 3), and a secondary term given by $\operatorname{def}(\varphi_{r})$. Indeed, it is well known that under a reasonable probability model, a random linear map tends to have a small defect. Unfortunately, if ${\mathsf{C}}$ is a random code (say, uniform among codes of given $[n,k]$), it is not easy to control the distribution of the maps $\varphi_{r}$.
Still, in the special case $r=2$, [3] manages to give lower bounds, exponentially close to $1$, on the probability that $\operatorname{def}(\varphi_{2})=0$. One might wonder how this generalizes. Let us review some arguments that support such a generalization, but also impose some limitations on its validity.
Reformulated in our setting, the minimal resolution conjecture of [18] postulates that, over an infinite field, a Zariski-generic code has $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$, or equivalently, that $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$. However, two points require our attention:
@@
    This conjecture is known to be false in general [10].
@@
Concerning point 1, we will argue that the conjecture is still “true enough” for us. First, a nonzero defect might not be a problem in our Betti number estimates, as long as it remains small. Moreover, as noted in the introduction of [10], the conjecture has been proved for a large range of values of $n$ and $k$. In fact, although [10] provides an infinity of counterexamples, these remain limited to specific parameters, namely of the form $n=k+O(\sqrt{k})$. And indeed, perhaps the most valuable result for us is [16], which proves that the conjecture is true when $n$ is large enough with respect to $k$.
Concerning point 2, it is true that codes behave quite differently over a finite field and over an infinite field. For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not. This makes it desirable to investigate links between Betti numbers and distance properties of a code.
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&881&1371&1122&315&63&6&-&-&-\\ 2&-&-&1&18&132&645&1680&1876&1221&485&110&11\\ \end{array}$
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1392&1181&300&49&4&-&-&-\\ 2&-&-&4&39&191&630&1666&1874&1221&485&110&11\\ \end{array}$
Figure 7: some actual $[23,12]_{2}$-codes with $d=4$, $d^{\perp}=4$ (compare also with Figure 3)
@@
    For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})=A_{d^{\perp}}({\mathsf{C}}^{\perp})$.
@@
    Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$. Moreover, quite often (but not always) we have $\beta_{k-d,k+1-d}({\mathsf{C}})=A_{d}({\mathsf{C}})$.
@@
Conversely, for a random ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$. In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
As an illustration, Figure 13 in the Supplementary material presents statistics on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes.
@@
    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$
@@
    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$.
@@
Consider this Heuristic in the asymptotic regime. Setting $R=k/n$, we can take $d=\operatorname{d}_{GV}(q,n,k)\approx H_{q}^{-1}(1-R)n$ and $d^{\perp}=\operatorname{d}_{GV}(q,n,n-k)\approx H_{q}^{-1}(R)n$ the corresponding Gilbert-Varshamov distance. Then the condition in 1. translates as $H_{q}^{-1}(R)>R^{2}$, and the condition in 2. translates as $H_{q}^{-1}(1-R)>R(1-R)$, both of which are satisfied when $R$ is small enough. In particular for $q=2$, we find that 1. is satisfied for $R<0.141$ and 2. is satisfied for $R<0.277$.
5 The distinguisher, with and without shortening
General principles
Distinguishers for alternant or Goppa codes tend to work better when $t$ is small and $n$ is large. When benchmarking, it is thus common to:
  • first fix $q,m$, and then for $n=q^{m}$, find the largest distinguishable $t$
  • once such a $t$ is found, fix $q,m,t$, and find the smallest distinguishable $n$.
Also, distinguishers typically work by computing certain code invariants. We might have theoretical bounds on the values of these invariants, that are essential for an asymptotic analysis. However these bounds need not be tight. Hence for a given set of finite parameters, we can also adopt a more empirical approach: sample a certain number of codes, compute their invariants, and observe when the distinguisher “just works”.
@@ 5 The distinguisher, with and without shortening
On the other hand, for random codes of dimension $k=mt$, if Heuristic 1 applies, (86) gives that $\beta_{r-1,r}$ should be equal, or very close, to $\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$. Thus we expect to distinguish when this value is smaller than $\beta_{r-1,r}^{*}$, or equivalently when
$n\geq\left\lceil\frac{k(k+1)}{r}-\frac{\beta_{r-1,r}^{*}-1}{\binom{k-1}{r-2}}\right\rceil$ (88)
(provided $\beta_{r-1,r}^{*}>0$).
@@
As for shortening, we do not expect it to improve the distinguishability threshold. This technique is used only, but crucially, to reduce the overall complexity when parameters grow larger.
Now let us see how our syzygy distinguisher fares, first on the examples proposed in [6], and then on the Classic McEliece system.
Goppa codes with $q=4$, $m=4$, $t=4$, irreducible Goppa polynomial
Goppa codes with $q=2$, $m=6$, $t=3$, irreducible Goppa polynomial
What happens? It turns out the conditions in Heuristic 1 are not satisfied anymore to ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$. From Experimental fact 1, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$. For $r=7$ and $k=mt=18$ this gives $d\geq 13$. For smaller $d$ we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$. The distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ w.h.p., and experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably. Now, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
The Classic McEliece 348864 system
We consider a dual code ${\mathsf{C}}\in\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$ proper of dimension $k=mt=768$. Theorem 3.1 gives $e=5$, $r_{\max}({\mathsf{C}})\geq 321$, which Corollary 4 improves to $\widehat{e}=3$, $r_{\max}({\mathsf{C}})\geq 427$.
For $n=3488$ and $k=768$ we have $\left\lfloor\frac{k(k+1)}{n}\right\rfloor=169$. Moreover, $d=\operatorname{d}_{GV}(2,n,k)=810$, so $k+1-d<0$. By Heuristic 1 we expect to distinguish at $\beta_{r-1,r}$ with $r=170$, with a large margin of safety. For this, Algorithm 1 computes the left kernel of the matrices $\mathbf{M}_{i}$ for $i\leq r$, where $\mathbf{M}_{i}$ has size $k\beta_{i-2,i-1}\times\binom{k+1}{2}\beta_{i-3,i-2}$. Using $\operatorname{ind}(\varphi_{i})=\left(\frac{k(k+1)}{i}-n\right)\binom{k-1}{i-2}$ as a lower bound for these $\beta_{i-1,i}$, we find that the maximal dimension of these matrices is at least $2^{596}$. Using $\omega\approx 2.372$ for the exponent of linear algebra gives a complexity at least $2^{1414}$.
However we have plenty of space in the interval $170\leq r\leq 427$ to apply shortening. Set $s=377$, $n_{s}=n-s=3111$, $k_{s}=k-s=391$. If ${\mathsf{C}}_{s}$ is a $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,12,3488,64}$, Corollary 4 ensures $r_{\max}({\mathsf{C}}_{s})\geq 427-s=50$. On the other hand, we have $\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor=49$, and $d_{s}=\operatorname{d}_{GV}(2,n_{s},k_{s})=921$, so $k_{s}+1-d_{s}\ll 0$. Thus we expect to distinguish at $\beta_{r_{s}-1,r_{s}}$ with $r_{s}=50$. Now, for $i\leq r_{s}$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}$. As $d_{s}^{\perp}=\operatorname{d}_{GV}(2,n_{s},n_{s}-k_{s})=55>r_{s}$, these $\beta_{i-1,i}$ can be estimated as $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$. We then find that the maximal dimension of these matrices is around $2^{223}$, and the complexity around $2^{529}$.
@@
Fix a base field cardinality $q$, for instance $q=2$, and a (dual) rate $R$. In [4] it is suggested to work with a primal code of rate between $0.7$ and $0.8$, so passing to the dual gives $0.2\leq R\leq 0.3$. For $n\to\infty$ we set:
@@
Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\perp}$, $\operatorname{{\mathsf{Gop}}}^{\operatorname{sqfr},\perp}$, or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes. Assume we have a lower bound of the form
$r_{\max}({\mathsf{C}}_{s})\geq k-s^{*}-s$ (91)
with $s^{*}=s^{*}(\mathcal{T}_{q,m,n,t})=o(n)$, valid for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of ${\mathsf{C}}$. Let then $r=o(n)$ be an integer such that $\frac{(s^{*}+r)(s^{*}+r+1)}{n-k+s^{*}+r}<r$. Now let ${\mathsf{C}}$ be a $[n,k]_{q}$-code. Then computing
$\beta_{r-1,r}({\mathsf{C}}_{k-s^{*}-r}),$ (92)
where ${\mathsf{C}}_{k-s^{*}-r}$ is a $(k-s^{*}-r)$-shortening of ${\mathsf{C}}$, allows to distinguish with high probability whether ${\mathsf{C}}$ is a random code ($\beta_{r-1,r}=0$) or an element of $\mathcal{T}_{q,m,n,t}$ ($\beta_{r-1,r}>0$), by doing at most $r$ kernel computations of linear maps between vector spaces whose maximal dimension can be estimated as $(s^{*}+r)^{3}\binom{s^{*}+r}{r-3}$.
In fact it is possible to take $s^{*}\approx R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n$ and $r\approx\frac{R^{2}}{1-R}\left(\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}\right)^{2}n$, which gives for the total complexity of the process an estimated upper bound of
$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}.$ (93)
Observe [5, §3.4] that the security parameter in the Classic McEliece system, based on the complexity of generic decoding algorithms, is linear in $t\propto n/\log(n)$. Then in (93) we have $\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$, which means that our complexity is subexponential in the security parameter.

### v5

@@ 4 Regularity $2$ and the small defect heuristic
Generating series reformulation of [10, Cor. 1.10].
Definition 5 ([26, Def. 1.5 & Th. 2.35])
@@
Definition 6 (cf. [10, §4A], after [23, Lect. 14])
@@
$\operatorname{ind}(\varphi_{r})=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ (83)
@@
Regularity $2$ helps in computing the full Betti diagram of codes. See Propositions 10 and 11 in the Supplementary material for more.
@@
By (78) and Theorem 4.1, for any $[n,k]$-code ${\mathsf{C}}$ of regularity $2$ and for any $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}.$ (85)
We might interpret this lower bound as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$. Then, when ${\mathsf{C}}$ is random, it is quite natural to expect that no other syzygy will appear, or equivalently that (85) will be an equality, i.e. $\operatorname{def}(\varphi_{r})=0$.
This is compatible with results such as those in [6] and the references within, that give estimates on the distribution of the rank, hence the defect, of random linear maps, under various probability laws. Unfortunately, even if the probability distribution of ${\mathsf{C}}$ is nice (say, uniform among codes of given $[n,k]$), it is not easy to control the distribution of $\varphi_{r}$, so we will not be able to give proofs. Moreover it could be that the iterative algebraic process in the construction of $\varphi_{r}$ would lead to some unexpected constraints on its rank. And indeed, in this section we will identify some parameter ranges for which the defect cannot be small; but conversely, we will also give arguments that support the validity of this small defect heuristic at least under some proper conditions.
A first argument in support of the heuristic is that it is unconditionally true when $r=2$. Indeed, in this special case, [3] manages to give lower bounds, exponentially close to $1$, on the probability that $\operatorname{def}(\varphi_{2})=0$.
Another argument is the minimal resolution conjecture of [19]. It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$ or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$. However, two points require our attention:
@@
    This conjecture is now known to be false in general [11].
@@
Concerning point 1, we will argue that the conjecture is still “true enough” for our use. First, a nonzero defect might not be a problem in our Betti number estimates, as long as it remains small. Moreover, as noted in the introduction of [11], the conjecture has been proved for a large range of values of $n$ and $k$. In fact, although [11] provides an infinity of counterexamples, these remain limited to very specific parameters, namely of the form $n=k+O(\sqrt{k})$. And indeed, perhaps the most valuable result for us is [17], which proves that the conjecture is true when $n$ is large enough with respect to $k$.
Concerning point 2, clearly, codes behave quite differently over a finite field and over an infinite field. For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not. This makes it desirable to investigate links between Betti numbers and distance properties of a code.
@@
    For all $r\geq d^{\perp}$ we have $\beta_{r-2,r}({\mathsf{C}})>0$.
    Moreover if $d^{\perp}\leq\frac{k(k+1)}{n}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
@@
    Dually, for all $r\leq k+1-d$ we have $\beta_{r-1,r}({\mathsf{C}})>0$.
    Moreover if $k+1-d\geq\frac{k(k+1)}{n}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
@@
Experimental fact 2
Conversely, for a random code ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$.
In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
Figures 13-14 in the Supplementary material illustrate various aspects of these Experimental facts.
@@
    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$
@@
    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$.
@@
Consider this Heuristic in the asymptotic regime. Setting $R=k/n$, we can take $d=\operatorname{d}_{GV}(q,n,k)\approx H_{q}^{-1}(1-R)n$ and $d^{\perp}=\operatorname{d}_{GV}(q,n,n-k)\approx H_{q}^{-1}(R)n$ the corresponding Gilbert-Varshamov distances, where $H_{q}$ is the $q$-ary entropy function. Then the condition in 1. translates as $H_{q}^{-1}(1-R)>R(1-R)$, and the condition in 2. translates as $H_{q}^{-1}(R)>R^{2}$, both of which are satisfied when $R$ is small enough. In particular for $q=2$, we find that 1. is satisfied for $R<0.277$ and 2. is satisfied for $R<0.141$.
5 The distinguisher
Basic version (without shortening)
Distinguishers typically work by computing certain code invariants. We might have theoretical bounds on the values of these invariants, that are essential for an asymptotic analysis. However these bounds need not be tight. Hence for a given set of finite parameters, we can also adopt a more empirical approach: sample a certain number of codes, compute their invariants, and observe when the distinguisher “just works”.
@@ 5 The distinguisher
For a given $r$ with $\beta_{r-1,r}^{*}>0$, our distinguisher then works as follows. Taking as input a (dual) code ${\mathsf{C}}$ of dimension $k=mt$, compute $\beta_{r-1,r}({\mathsf{C}})$ with Algorithm 1. Then if $\beta_{r-1,r}({\mathsf{C}})\geq\beta_{r-1,r}^{*}$ declare ${\mathsf{C}}$ is of the special type in question. Otherwise declare ${\mathsf{C}}$ is random.
Observe that for random codes of dimension $k$, if the conditions in Heuristic 1 are satisfied, we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$. Thus the distinguisher succeeds when this value is smaller than $\beta_{r-1,r}^{*}$, or equivalently when
$n\geq\left\lceil\frac{k(k+1)}{r}-\frac{\beta_{r-1,r}^{*}-1}{\binom{k-1}{r-2}}\right\rceil.$ (86)
@@
We compared our new distinguisher (in this basic version) with the one from [7], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$. Results can be found at the end of the Supplementary material. In both cases, we improve the distinguishability threshold. Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (86) (with $r=r_{\max}^{*}=4$), i.e. the conditions in Heuristic 1 are always met. On the other hand, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$, the conditions in Heuristic 1 are the limiting factor (but still improving over [7]).
A general description of the set of distinguishable parameters is complicated, as it has to encompass both (86) and the conditions in the heuristic. However, a very interesting phenomenon will appear in the asymptotic analysis: all these constraints will be automatically satisfied, i.e. asymptotically all parameters will become distinguishable.
Full version (with shortening)
We can also use shortened dual codes as input in our distinguisher. In general, shortening $s$ times a $[n,k]$-code ${\mathsf{C}}$ produces a $[n_{s},k_{s}]$-code ${\mathsf{C}}_{s}$ with $n_{s}=n-s$, $k_{s}=k-s$. Having smaller parameters will decrease the complexity. Moreover the (dual) rate $R_{s}=k_{s}/n_{s}$ also decreases when we shorten; as observed in Remark 2, this will help satisfying the conditions in Heuristic 1.
Example 1
Consider the class $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ of dual binary Goppa codes with $m=10$, $n=1024$, $t=10$, irreducible Goppa polynomial. Without shortening, these codes are out of reach of our distinguisher: we would have to run Algorithm 1 up to $r=10$ and do linear algebra in dimension roughly $10^{20}$, which is infeasible. On the other hand, we can effectively distinguish after shortening $40$ times: we experimentally find that $40$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ have $\beta_{3,4}=30$, while $\beta_{3,4}=0$ for random codes with the same parameters. Computational cost (several days each, on a workstation with several hundreds of GB of RAM) prevented the author to run more than a few tests, but no deviation was observed. In particular $\beta_{3,4}=0$ consistently in the random case confirms Heuristic 1 for these parameters.
Proposition 8
Fix a class $\mathcal{T}_{q,m,n,t}$ of dual alternant or Goppa codes with some given parameters $q,m,n,t$, and assume that for some integer $r^{*}$ we have a lower bound of the form
$r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ (87)
for all ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$ and all $s$-shortened subcodes ${\mathsf{C}}_{s}$ of ${\mathsf{C}}$.
Set $k=mt$, and let ${\mathsf{C}}$ be either a random $[n,k]_{q}$-code, or an element of $\mathcal{T}_{q,m,n,t}$.
Choose an integer $s$ such that
$\frac{k_{s}(k_{s}+1)}{n_{s}}<r^{*}-s$ (88)
where $n_{s}=n-s$ and $k_{s}=k-s$.
  1. 1.
    Let $d_{s}\approx\operatorname{d}_{GV}(q,n_{s},k_{s})$ be the typical minimum distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$. Then, under part 1. of Heuristic 1, computing
    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s}),$ (89)
    where ${\mathsf{C}}_{s}$ is any $s$-shortened subcode of ${\mathsf{C}}$, allows to distinguish whether ${\mathsf{C}}$ is a random code (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ w.h.p.) or an element of $\mathcal{T}_{q,m,n,t}$ (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$).
  2. 2.
    Moreover, let $d_{s}^{\perp}\approx\operatorname{d}_{GV}(q,n_{s},n_{s}-k_{s})$ be the typical dual distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$. Then, under Heuristic 1, for large enough parameters the complexity of this distinguisher is dominated by the sum
    $\kappa=\sum_{i\leq\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1}\max\left(k_{s}\operatorname{ind}(\varphi_{i-1}),\binom{k_{s}+1}{2}\operatorname{ind}(\varphi_{i-2})\right)^{\omega}$ (90)
    where $\operatorname{ind}(\varphi_{i})=\left(\frac{k_{s}(k_{s}+1)}{i}-n_{s}\right)\binom{k_{s}-1}{i-2}$ and $\omega\approx 2.372$ is the exponent of linear algebra.
Proof
If ${\mathsf{C}}$ is random, so is ${\mathsf{C}}_{s}$. Then, as $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 1. of Heuristic 1 we indeed expect $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ for $r^{*}-s>\frac{k_{s}(k_{s}+1)}{n_{s}}$. On the other hand, for ${\mathsf{C}}$ in $\mathcal{T}_{q,m,n,t}$, the lower bound $r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ precisely means $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$.
Now using Algorithm 1 we obtain $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $2\leq i\leq r^{*}-s$. The contributions for $i<4$ are negligible. For $i\geq 4$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}({\mathsf{C}}_{s})\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{s})$ and the contributions for $i>\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1$ are negligible because of part 1. Then, as $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 2. of Heuristic 1 we expect $\beta_{i-1,i}({\mathsf{C}}_{s})=\operatorname{ind}(\varphi_{i})$ for $2\leq i\leq\frac{k_{s}(k_{s}+1)}{n_{s}}$ for random ${\mathsf{C}}$; and we can assume this holds also for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, otherwise computing this dimension readily provides a distinguisher with lower complexity. From this we conclude.
Example 2
The following table shows how the distinguisher from Proposition 8 applies to Classic McEliece parameters. We take $r^{*}=\widehat{f}$ as given by Corollary 4, and then choose the maximum possible shortening order $s$ satisfying (88).
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr(n,m,t)&(3488,12,64)&(4608,13,96)&(6688,13,128)&(6960,13,119)&(8192,13,128)\\ \hline\cr r^{*}&427&683&939&867&939\\ \hline\cr s&377&568&816&769&848\\ \hline\cr r^{*}-s&50&115&123&98&91\\ \hline\cr[n_{s},k_{s}]&[3111,391]&[4040,680]&[5872,848]&[6191,778]&[7344,816]\\ \hline\cr\frac{k_{s}(k_{s}+1)}{{n_{s}}_{\phantom{0\!\!}}}&49.27&114.62&122.61&97.89&90.78\\ \hline\cr\operatorname{d}_{GV}&921&1069&1650&1828&2256\\ \hline\cr\operatorname{d}_{GV}^{\perp}&55&62&122&108&110\\ \hline\cr\kappa&{2^{528}}^{\phantom{i\!}}&(2^{1080})&(2^{1224})&2^{1030}&2^{997}\\ \hline\cr\end{array}$
We observe that the condition $\operatorname{d}_{GV}(n_{s},k_{s})>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$ is always satisfied with a large margin, so we could choose a smaller $s$ and still be able to distinguish. Concerning the complexity estimate $\kappa$, it relies on the condition $\operatorname{d}_{GV}(n_{s},n_{s}-k_{s})>\frac{k_{s}(k_{s}+1)}{n_{s}}$. We see this condition is satisfied for the parameter sets $(3488,12,64)$, $(6960,13,119)$, $(8192,13,128)$, but for $(4608,13,96)$ and $(6688,13,128)$ it is not (then, it would still be possible to give a complexity estimate, using Lemma 5 instead of Heuristic 1). These complexities improve those from [7], although they remain practically unreacheable and well beyond security levels.
@@
Fix a base field cardinality $q$, for instance $q=2$, and a (dual) rate $R$. In [4] it is suggested to take a primal code of rate between $0.7$ and $0.8$, so passing to the dual gives $0.2\leq R\leq 0.3$. However here we allow any $R$. Then for $n\to\infty$ set:
@@
A key observation then is that the lower bound $f$ on $r_{\max}$ of dual alternant codes in Theorem 3.1, or equivalently, the number $f$ of columns of the matrix $\mathbf{\Phi}$ in (65), is very close to $k$:
Lemma 11
We have
Proof
Direct consequence of:
  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$.
In the case of binary Goppa codes, the $\widehat{f}$ from Corollary 4 improves the $f$ from Theorem 3.1. However one could show that both have the same asymptotics. So, in order to distinguish these codes from random codes, we will not care about the extra structure: treating Goppa codes as alternant codes will suffice. (Still, this could be used to distinguish Goppa from “general” alternant codes.)
Now, under Heuristic 1, we have:
@@
Asymptotically, $q$-ary alternant (including Goppa) codes of dual rate $R$ can be distinguished from random codes, with complexity at most
$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}$ (92)
where $\omega\approx 2.372$ is the exponent of linear algebra.